Please show details of solution. Thanks

Ann purchased a property for $1,000,000. She bought the property at a 6.00% cap rate.

She finances the purchase with an Interest Only senior loan at 50% LTV at an interest rate of 4.00%.

She also decides to get subordinate / mezzanine financing for 20% of the capital stack (from 50%-70% LTV) at 8.00% interest only.

What is the weighted average cost of Ann’s debt?
Answer #1

1. Value of Interest only Loan = Property Value * 50% = 1000000 * 50% = 500000

2. Value of Subordinate Financing = Property Value * 20% = 1000000 * 20% = 200000

3. weighted average cost of Ann’s debt = Weight of Loan 1 * Cost of Loan 1 + Weight of Loan 2 * Cost of Loan 2

weighted average cost of Ann’s debt = 5000000/7000000 * 4% + 2000000/7000000 * 8%

weighted average cost of Ann’s debt = 5.14%

Please dont forget to upvote